OpenStudy (anonymous):

How do I solve 4(2x + 1)- 3 = -3(x - 4)?

OpenStudy (anonymous):

You have to distribute out what's outside the parentheses first. Do you know how to do that?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Distributive Preoperty: \(a(b + c) = ab + ac\)

OpenStudy (anonymous):

Oh wait yes.

OpenStudy (anonymous):

That's basically means that the number outside of the parentheses you have to multiply it to everything inside the parentheses

OpenStudy (anonymous):

I got 8x+4-3=-3x+12+12, is it right?

OpenStudy (anonymous):

8x+4-3=-3x+12 You added an extra 12 on the right hand side

OpenStudy (anonymous):

But other than that you are right. Think you can handle it from there?

OpenStudy (anonymous):

Oh oops, lol.

OpenStudy (anonymous):

Now what do I do?

OpenStudy (anonymous):

You want all the x's to be on one side and all the numbers on the other side. Let's start by adding 3x to both sides. This will get rid of the -3x from the right hand side. 8x+4-3=-3x+12 8x+3x+4-3=-3x+3x+12 8x+3x+4-3=12 11x+4-3=12

OpenStudy (anonymous):

From there you can simplify it even further by doing this 4-3?

OpenStudy (anonymous):

Would the answer be 1, right?

OpenStudy (anonymous):

Yeah you are right man Good job!

OpenStudy (anonymous):



OpenStudy (anonymous):

Thank you =)